ISC Meeting on 2016/1/25

14:00 - 15:00 on SeeVogh

Participants: Arai, H. Yamamoto, Michimura, Somiya, Aso, Hirose, Miyakawa, Kokeyama

  1. Schedule for iKAGRA (Aso)
    • Installation of the analog electronics
    • Installation of the detection optics
      • Above two will be done in the week of 2/15
    • Initial alignment procedure
      • Michimura-kun will make a document
    • Digital system integration and lock acquisition
      • Need to test the system beforehand: Will consult with Kokeyama-san
      • Aso will try to gather man power (mostly students)
  2. Electronics Preparation Status (Michimura, Aso)
    • RF PD (Aso, Doi)
      • Expected delivery 2/15
    • Cable assembly
      • RF flexible cables with N connectors: ordered
      • Spare RF flexible cables with SMA connectors: waiting for Ohashi's approval
      • Diff-to-single cables: to be made at the time of installation
      • we have purchased all the other ISC cables
  3. Magnet noise
    • Magnetometer requirement(Michimura)
      • Magnetometer requirement: According to Shimoda report (JGW-T1504459), the magnetic field fluctuation of 0.5 pT/rtHz @ 10 Hz creates the displacement noise of 1e-20 m/rtHz @ 10 Hz for bKAGRA ITM/ETM. If we keep the magnet size as current design (2mm dia, 2mm thick), the magnetic field fluctuation should be smaller than 7 pT/rtHz @ 10 Hz, considering the displacement noise requirement is 1.5e-19 m/rtHz @ 10 Hz. So, 7 pT/rtHz @ 10 Hz will be the requirement for the magnetometer noise (Bartington Mag-03 meets this requirement). However, considering the magnetic field measurement done by Virgo (30 pT/rtHz @ 10 Hz with equipments installed), we might have to change the magnet size to even smaller one.
      • Measurement in LIGO: About 10pT/rtHz in front of an electronics rack
      • MIF requirement: we want to see 7pT
  4. FFT simulation (H. Yamamoto, Hirose)
    • ETM polishing will take until the end of March. -> Simulations with measured transmission maps can be performed only after April.

    • The contract for the ITM polish cannot wait for the simulation result
      • We will target 6nm RMS of equivalent optical path fluctuation for ITMs.
    • Recent mirrors from the US company have large absorption (~ 100ppm/cm)
    • One of the two samples from a Japanese company showed a low absorption (~ 30ppm/cm)
      • Will measure the transmission map after polishing for this sample
  5. Next meting

2016/2/16 14:00 - 15:00

KAGRA/subgroup/ifo/MIF/Minutes20160125 (last edited 2016-02-01 11:42:10 by YoichiAso)